Suppose u; v € V and ||u|| = ||v|| = 1 with < u; v > = 1: Prove that u = v


1
Expert's answer
2021-07-12T14:02:35-0400

<u,v>=||u||"\cdot" ||v||cos"\theta"

1=1"\cdot"1cos"\theta"

cos"\theta=1"

"\theta=0\degree"


So, the magnitudes of the vectors are equal, and the angle between them is 0"\degree". That is the vectors are equal.